The finished inside diameter of a piston ring is normally distributed with a mean of 10 centimeters and a standard deviation of 0.03 centimeter.

(a) What proportion of rings will have inside diameters exceeding 10.075 centimeters?

(b) What is the probability that a piston ring will have an inside diameter between 9.97 and 10:03 centimeters?

(c) Below what value of inside diameter will 15% of the piston rings fall?


1
Expert's answer
2021-07-06T07:24:30-0400

Let X represent that the finished inside diameter of a piston ring is normally distributed with the mean 10 centimeters and a standard deviation of 0.03 centimeters.

(a) The proportion of rings will have inside diameters exceeding 10.075 centimeters

"z=\frac{x- \mu}{\sigma}= \frac{10.075-10}{0.03}=2.5 \\

P(X>10.075) = P(Z>2.5) \\

= 1 - P(Z<2.5) \\

= 1 -0.9938 \\

= 0.0062 = 0.62 \; \%"

(b) The probability that a piston ring will have an inside diameter between 9.97 and 10:03 centimeters

"P(9.97<X<10.03) = P(X<10.03) -P(X<9.97) \\

=P(Z< \frac{10.03-10}{0.03}) -P(Z< \frac{9.97-10}{0.03}) \\

=P(Z< 1) -P(Z< -1) \\

= 0.8413 -0.1587 \\

= 0.6826"

(c) The value of inside diameter that will 15% of the piston rings fall

"P(X<x) = 0.15 \\

P(Z< \frac{x-10}{0.03}) = 0.15 \\

P(Z< -1.036) = 0.15 \\

\frac{x-10}{0.03}= -1.036 \\

x-10 = - 1.036 \times 0.03 \\

x= -1.036 \times 0.03 + 10 \\

x = 9.9689"